# A low complexity contextual stacked ensemble-learning approach for pedestrian intent prediction

### v1

@@ V Conclusion
This work examines the critical problem of pedestrian intent prediction (PIP), which pauses safety concerns for vulnerable road users (VRU) in the advent of driver-assisted and driver-less vehicles. This gained ground in ensuring VRUs safety in the presence of driver-less and driver-assisted vehicles directly affects the uptake of active travel in urban areas and thus, accelerates the transition toward sustainable transportation and mobility solutions.
Given rising concerns about the carbon footprint of artificial intelligence, we propose a Contextual Stacked Ensemble-learning (CSE) approach to analyze visual data taken from an ego-vehicle with $99.7\%$ reduction in computational complexity compared to the state-of-the-art and competitive performance with Pedgraph+ and Global PCPA with comparable performance. Our method benefits from three pillars: the first relates to data compression using skeleton-ization, the second to contextualization based on in-depth analysis of attributes inferred from visual data, and the third to trajectory analysis of VRUs. An extension of this work might examine how much information the network can contain and evaluate our ensemble technique presented in this paper. In addition, the training of the three different output heads is done separately instead of simultaneously. Training in a multi-task formulation on multiple GPUs could be explored in the future.

### v2

@@ V Conclusion
This work examines the critical problem of pedestrian intent prediction (PIP), which raises safety concerns for vulnerable road users (VRU) in the advent of driver-assisted and driver-less vehicles. Ensuring the safety of VRUs when these types of vehicles are present is crucial, as it impacts the adoption of active travel in urban areas. An increase in the uptake of active travel accelerates the transition to sustainable mobility and transportation solutions.
With growing concerns about the carbon footprint of artificial intelligence, we propose a Contextual Stacked Ensemble-learning (CSE) approach to analyze visual data collected from ego-vehicles. Our method achieves a $99.97\%$ reduction in computational complexity compared to state-of-the-art techniques while demonstrating competitive performance with Pedgraph+ and Global PCPA. The effectiveness of our approach is built on three main pillars: the first focuses on data compression through skeleton-ization; the second emphasizes contextualization based on a thorough analysis of attributes inferred from visual data; and the third involves trajectory analysis of VRUs.
An extension of this work might examine how much information the network can contain and evaluate our ensemble technique presented in this paper. Moreover, we argued that when ensemble learning can(not) boost performance and how to evaluate input data and model diversity. Finally, the training of the three different output heads is done separately instead of simultaneously. Training in a multi-task formulation on multiple GPUs could be explored in the future.